Numbers 1 and 2 please
• Giren f(x, y, z) = ye + x1nz Find E, ly, la, tra, Exy 2 Given w= xy + y +XZ. X= s. cost y=s. sint z =t Find dw and dw dt
Answer #1

- 0 Given f(x, y, z) = yeatainz a Fird foc, fy, focol, fyy, forg solution f(41.4,2) = yes + onlin2 E fx = 2 (y e8t 2l22) Assyfog bolig (yellty/m23] fony - Cenyt silmz)] by len yt inz] e9,170 & Given w = nyty z +22 7= 5. copt , y = $. Sin #, z= Gnot fSolution w=ayt gz + XZ put all valuepf 7, 9,2 Was sin # )(scopP+#ssin t) + scoot dW = f ( 8 sinf cooft Assinat of cool] ds do